OpenStudy (anonymous):

write in simplest form sin inverse(3x-4xcube)

OpenStudy (anonymous):

Just make a little substitution here... Put x = sinx here..

OpenStudy (anonymous):

\[\sin^{-1}(3six - 4\sin^3x) \implies \sin^{-1}(\sin3x) = 3x\] \[= 3\sin^{-1}x\]

OpenStudy (anonymous):

itna hi ?

OpenStudy (anonymous):

Tumhe isse bhi simplest form chaahiye??

OpenStudy (anonymous):

nhi thanku

OpenStudy (anonymous):

Welcome dear..